the sum of two consecutive integers is 138. find the integers.
Answer by TutorDelphia(193)

Two consecutive integers will not add up to be an even number since an odd number plus an even number always results in an odd (try it odd 3+4=7; 12+13=25)
I'm guessing you mean the sum of three consecutive integers
lets name our first number n. Our next number in line will be one more than that...so n+1 followed by n+2
all these add up to be 138
so we get
n+(n+1)+(n+2)=138 (the parenthesis don't do anything since this is all addition but help to separate the three consecutive integers)
combine like terms
3n+3=138
subtract 3 from both sides
3n=135
divide both sides by 3
n=45
n+1=46
n+2=47
and 45+46+47=135
so those are your three numbers